Need Help with this problem. Due tonight and don't understand it.
2)Is (sqrt x)^2 an identity (true for all nonnegative values of x)?
Answer:
Explain your answer in this space.

Answer by stanbon(75887)

Yes, true for all non-negative values of "x".
sqrt(anything) is always non-negative, if it is a Real number.
What about sqrt(x^2)? Does it equal "x"?
sqrt(-2)^2 = sqrt4 =2
sqrt(0)^2 = sqrt0 = 0
sqrt(+2)^2= sqrt4= 2

So, sqrt(x^2) is not always "x".
The sqrt (anything) is no-negative, if it is a Real number.
Cheers,
Stan H.